# On Computability of Computable Problems

### v1

@@ 3 Normalization of $f(Z,T)$ Transform
Suppose there is problem $P$ that is computable for input $\psi$, and $\psi\in\Psi$. We apply $f(Z,T)$ transform on $P(\psi)$ and get $P(t)(\psi)$.Normalization of the transformed problem $P(t)(\psi)$ is defined in following way.
@@
Because $P(\psi)$ is computable problem and $0<T_{P}\leq\infty$ therefore $0\leq f_{n}(P,T)<\infty$.The problem $P(t)(\psi)$ is computable if $10\leq f_{n}(P,T)<$ 1 because $P(t)(\psi)$ halts before Stipulation machine rewrites input. If $N_{P}T_{t}>T_{P}$ then $f_{n}(P,T)>1$ i.e., input of $P(t)$ varies before $P(t)(\psi)$ gets computed and $P(t)$ halts. Given that $\psi\in\Psi$ and $\Psi=\{\psi_{1},\text{ }\psi_{2},\text{ }\psi_{3},\text{ }\ldots,\text{ }\psi_{n}\}$, and according to postulate V, no two consecutive inputs of $P(t)$ are identical.If $f_{n}(P,T)>1$ then Stipulation machine rewrites input on $P(t)$ before $P(t)$ decides and halts on input $\psi$.Consequently, $P(t)$ carries out computation on some parts of each recurring input in some order.Those some parts and some order make arbitrary string $\psi_{f}$ of some arbitrary length, and composition of $\psi_{f}$ depends on $T_{P}$.The string $\psi_{f}$ may not be member of set $\Psi$.Therefore, if $f_{n}(P,T)>1$ then $P(t)$ is not computable when input is $(\psi)$, but if $10\leq f_{n}(P,T)<$ 1 then $P(t)$ is computable when input is $(\psi)$.Normalization line for $P(t)(\psi)$ is shown in Fig.3.The set $\Psi$ can have input strings of any length, and this important aspect (infinite cardinality of set $\Psi$) is discussed in Section 5.
@@
Point of inflection and point of convergence are two important points on normalization line. If $f_{n}(P,T)=0$ then transformed problem $P(t)(\psi)$ exist at point of convergence, and $f(P,T)=f(P)$ when $f_{n}(P,T)=0$.If $f_{n}(P,T)=1$ then transformed problem $P(t)(\psi)$ exist at point of inflection.The point where $f_{n}(P,T)=1$ is very critical point for computability of problem $P(t)(\psi)$.A small change can make $P(t)(\psi)$ move in-between computable-uncomputable when $f_{n}(P,T)$ exists around point of inflection.If $P(\psi)$ requires $N_{P}$ steps and $P(t)(\psi)$ exists around point of inflection, then slight change in $N_{P}$ (e.g., $N_{P}-n$ or $N_{P}+n$ for some $n\geq 1$) or $T_{P}$ can change the computability of $P(t)(\psi)$.We can also represent normalization line of $P(t)(\psi)$ in following way.
@@
The $f(Z,T)$ transformed problems that exist just before or immediately after point of inflection are very helpful in analysing the computational characteristics of different variants of the Turing machines at infinitesimal scale.If cardinality of $f(Z,T)$ transformed problems is higher than cardinality of Turing machines (defined over finite set of states) then there must exist $f(Z,T)$ transformed problems that are uncomputable.Cardinality of $f(Z,T)$ transformed problems is discussed in Section 5.
4 Computation of $f(Z,T)$ Transformed Problems by variants of Turing machines
[…]
Figure 4: Computability of $P(t)(\psi)$ through variants of Turing machines
4.1 $k$-tape Turing machines for $P(t)(\psi)$ problems
If there is a string $m$ of length $n$ then $1$-tape Turing machine requires $O(n^{2})$ number of steps to decide whether $m$ is palindrome or not, and $2$-tape Turing machine requires $O(n)$ steps to make this decision about $m$.The problem $P$ is to decide twice whether the given input string is palindrome, if both answers are yes then accept, if both answers are no then reject; and if both answers are not same then start evaluating again and continue until we have two consecutive same answers.For a given input $m$, we transform $P(m)$ into $P(t)(m)$, and then adjust $T_{P}$ in such a way that $f_{n}(P,T)<1$ for $2$-tape Turing machines and $f_{n}(P,T)$ of $P(t)(m)$ exists just before point of inflection.Single tape and $2$-tape Turing machines are available, and both can recognize $P(m)$.Now, for this $P(t)(m)$ and $f_{n}(P,T)$, replace $2$-tape Turing machine connected to Stipulation machine with $1$-tape Turing machine.The problem $P(t)(m)$ is decidable for $2$-tape Turing machine because $f_{n}(P,T)<1$, but it is not decidable by $1$-tape Turing machine because $N_{P}$ is higher for $1$-tape Turing machine (as it requires $O(n^{2})$ steps) and consequently $f_{n}(P,T)>1$.The transformed problem $P(t)(m)$ moves from being decidable (i.e., $f_{n}(P,T)<1$) for $2$-tape Turing machine to undecidable (i.e., $f_{n}(P,T)>1$) for $1$-tape Turing machine.If there is a $k$-tape Turing machine that can decide $P(m)$ then, only for $k\geq 2$, it can decide this $P(t)(m)$ with given $f_{n}(P,T)$.
Suppose there is a set $\Pi$ that consists of $k$ number of binary strings, and each string has some finite length $n$ for some $n\geq 10$.The problem $P_{1}$ is to decide thrice that the binary strings of $\Pi$ are identical, if all three answers are yes then accept, if all three answers are no then reject; and if all answers are not same then start evaluating again and continue until we have three consecutive same answers.Suppose $P_{1}$ is decidable for a given $\Pi$.We transform $P_{1}(\Pi)$ into $P_{1}(t)(\Pi)$, and then adjust $T_{P_{1}}$ in such a way that $f_{n}(P_{1},T)<1$ for $k$-tape Turing machines, and $f_{n}(P_{1},T)$ exists just before point of inflection for $P_{1}(t)(\Pi)$.Now, for this $P_{1}(t)(\Pi)$ and $f_{n}(P_{1},T)$, replace $k$-tape Turing machine connected to Stipulation machine with $k-1$ tape Turing machine. The problem $P_{1}(t)(\Pi)$ is decidable by $k$-tape Turing machine but it is not decidable by $k-1$ tape Turing machine because $N_{P_{1}}$ will increase for $k-1$ tape Turing machine and consequently $f_{n}(P_{1},T)>1$.The problem $P_{1}(t)(\Pi)$ moves from being decidable (i.e., $f_{n}(P_{1},T)<1$) for $k$-tape Turing machine to undecidable (i.e., $f_{n}(P_{1},T)>1$) for $k-1$ tape Turing machine.Similarly, there exists a transformed problem that is computable for $k+1$ tape Turing machine but not computable for $k$ tape Turing machine.
We can construct a computable problem $P$ that, for a given input $\psi$, takes slightly more number of steps in its computation if we decrease the number of tapes.This slight increment in the number of computational steps provides opportunity to apply $f(Z,T)$ transform and get transformed problem $P(t)(\psi)$, we adjust its variation time $T_{P}$ such that $f_{n}(P,T)$ exist around point of inflection for $P(t)(\psi)$, and then present $P(t)(\psi)$ to different variants of Turing machines for computation.Depending on the variant, the problem $P(t)(\psi)$ can be computable or uncomputable, because $f_{n}(P,T)$ for $P(t)(\psi)$ is different for different variants.
If $T(k,2)$ represents $k$-tape Turing machine defined over binary input, then there exist problem $P(t)(\psi)$ such that Turing machines can be enlisted according to their computational power in following way for $f(Z,T)$ transformed problems:
$T(1,2)\subsetneq T(2,2)\subsetneq T(3,2)\subsetneq\ldots\subsetneq T(k,2)\subsetneq T(k+1,2)\ldots$ (31)
We can state that there exists a problem $P(t)(\psi)$ such that $T(k,2)\neq T(k+1,2)$ because $f_{n}(P,T)>1$ for $T(k,2)$ and $f_{n}(P,T)<1$ for $T(k+1,2)$
For a given problem $P(t)(\psi)$, if $f_{n}(P,T)=0$ because $T_{P}=\infty$ then $T(1,2)=T(k,2)$ for all finite $k$, i.e., if input of problem $P(t)$ does not vary at all (because $T_{P}=\infty$) then single tape Turing machine is computationally as powerful as any other Turing machine defined over finite number of tapes.To put it in another way, if problem $P(t)(\psi)$ exist at point of convergence then all Turing machines (defined over finite number of tapes) converge to equal computational power.
4.2 $w$-alphabet Turing machines for $P(t)(\psi)$ problems
As stated earlier in linear speedup theorem that increase in alphabet size results in tape compression, and that consequently decreases the number of steps required in a computation.Any change in number of steps required in a computation changes the normalization value of $f(Z,T)$ transformed problem, and a change in the normalization value of a transformed problem can cause outcome of computability for the transformed problem to change.
We need unary string of length $2^{x}$ to represent binary string of length $x$.Turing machines defined over unary input can have better time complexity only if the time complexity is measured in term of unary input length, but if time complexity is measured in term of total number of steps then Turing machines defined over unary input take more number of steps to compute a problem than that of Turing machines defined over binary or higher numeral systems.
The problem $P$ is to decide whether the given input string $m$ is palindrome or not, if both answers are yes then accept, if both answers are no then reject; and if both answers are not same then start evaluating again and continue until we have two consecutive same answers. For given input string $m$, we transform this problem $P(m)$ into $P(t)(m)$, and then adjust $T_{P}$ in such a way that $f_{n}(P,T)<1$ for a Turing machine with input alphabet $w=\{0,1\}$, and $f_{n}(P,T)$ exists just before point of inflection for $P(t)(m)$. If $N_{P}$ represents number of steps required to compute $P(m)$, then decrease in $|w|$ will increase $N_{P}$ and normalization value. There exist $m$ such that $P(t)(m)$ is computable if $|w|>1$ because $f_{n}(P,T)<1$, and $P(t)(m)$ not computable if $|w|=1$ because $f_{n}(P,T)>1$. Similarly, we can construct and show the existence of $f(Z,T)$ transformed problem that is computable if $|w|\geq g+1$ and it is not computable if $|w|=g$ for any positive integer $g$.
We can construct a computable problem $P$ that, for a given input $\psi$, takes slightly fewer number of steps in its computation if we increase alphabet size. This slight decrement in the number of computational steps provides opportunity to apply $f(Z,T)$ transform and get transformed problem $P(t)(\psi)$, we adjust its variation time $T_{P}$ such that $f_{n}(P,T)$ exist around point of inflection for $P(t)(\psi)$, and then present $P(t)(\psi)$ to different variants of Turing machines for computation.Depending on the variant of Turing machine, the problem $P(t)(\psi)$ can be computable or uncomputable, because $f_{n}(P,T)$ for $P(t)(\psi)$ is different for different variants.
If $T(1,w)$ represents single tape $|w|$-input alphabet Turing machine, then there exist problem $P(t)(\psi)$ such that Turing machines can be enlisted according to their computational power in following way for $f(Z,T)$ transformed problems:
$T(1,1)\subsetneq T(1,2)\subsetneq T(1,3)\subsetneq\ldots\subsetneq T(1,w)\subsetneq T(1,w+1)\ldots$ (32)
We can state that there exists a problem $P(t)(\psi)$ such that $T(1,w)\neq T(1,w+1)$ because $f_{n}(P,T)>1$ for $T(1,w)$ and $f_{n}(P,T)<1$ for $T(1,w+1)$.
For a given problem $P(t)(\psi)$, if $f_{n}(P,T)=0$ (because $T_{P}=\infty$) then $T(1,1)=T(1,w)$ for all finite set of $w$. To describe it in another way, if problem $P(t)(\psi)$ exist at point of convergence then all Turing machines, defined over finite set of input alphabet, converge to equal computational power.
4.3 $k$-tape and $w$-alphabet Turing machines for $P(t)(\psi)$ problems
A $k$-tape Turing machine defined over $w$ input alphabet is represented by $T(k,w)$. As mentioned in subsection 4.1 and 4.2, we can construct and show the existence of $f(Z,T)$ transformed problem $P(t)(\psi)$ such that $f_{n}(P,T)>1$ for $T(n,n)$ and $P(t)(\psi)$ is not computable, but the problem $P(t)(\psi)$ has $f_{n}(P,T)<1$ for $T(n+1,n+1)$ and $P(t)(\psi)$ is computable.
We can construct a computable problem $P$ that, for a given input $\psi$, takes slightly fewer number of steps in its computation if we increase alphabet size and number of tapes.This slight decrement in the number of computational steps provides opportunity to apply $f(Z,T)$ transform and get problem $P(t)(\psi)$, we adjust variation time $T_{P}$ in such a way that $f_{n}(P,T)$ exist around point of inflection for $P(t)(\psi)$, and present $P(t)(\psi)$ to different variants of Turing machines for computation.Depending on the variant, the problem $P(t)(\psi)$ can be computable or uncomputable, because $f_{n}(P,T)$ for $P(t)(\psi)$ is different for different variants.
There exist problem $P(t)(\psi)$ such that Turing machines $T(k,w)$ can be enlisted according to their computational power in following way for $f(Z,T)$ transformed problems:
$T(1,1)\subsetneq T(2,2)\subsetneq T(3,3)\subsetneq\ldots\subsetneq T(n,n)\subsetneq T(n+1,n+1)\ldots$ (33)
We can state that there exists a problem $P(t)(\psi)$ such that $T(n,n)\neq T(n+1,n+1)$ because $f_{n}(P,T)>1$ for $T(n,n)$ and $f_{n}(P,T)<1$ for $T(n+1,n+1)$
For a given problem $P(t)(\psi)$, if $f_{n}(P,T)=0$ (because $T_{P}=\infty$) then $T(1,1)=T(n,n)$ for all finite $n$, i.e., if input of problem $P(t)$ does not change at all (because $T_{P}=\infty$) then Turing machine defined over $k=1$ and $w=1$ is computationally as powerful as any Turing machine defined over finite number of tapes and finite set of input alphabets.
It can be stated that if problem $P(t)(\psi)$ exists at point of convergence then all Turing machines (defined over finite number of tapes and finite set of input alphabet) converge to equal computational power.All variants of Turing machine have equal computational power if $f(Z,T)$ transformed problems exist at point of convergence, i.e., $f_{n}(Z,T)=0$.
If there exist $f(Z,T)$ transformed problems that are uncomputable by Turing machines then cardinality of $f(Z,T)$ transformed problems must be higher than cardinality of Turing machines.
5 Cardinality of Problems Produced by $f(Z,T)$ Transform
@@
The $f(Z,T)$ transform is applied to only computable functions (or decidable languages) and it produces transformed problems.Suppose there is a decidable language $L_{C}$ and it has only finite number of strings, and each string in $L_{C}$ has length $n$, i.e., $l_{C}\in L_{C}$ and $|l_{C}|=n$ for all $l_{C}$.There exist a Turing machine $P_{C}$ that can recognize $L_{C}$, and $P_{C}(l_{C})$ is decidable for all $l_{C}$.We want to know the cardinality of $f(Z,T)$ transformed problems if we apply $f(Z,T)$ transform on $L_{C}$.We change $P_{C}$ into $P_{C}(t)$ and input is from $L_{C}$. If $0<T_{P_{C}}\leq\infty$ and $T_{P_{C}}\in\mathbb{Q}^{+}$ then $|T_{P_{C}}|=\aleph_{0}$.If $|T_{P_{C}}|=\aleph_{0}$ then $f(Z,T)$ transform produces infinite number of transformed problems in the form of $P_{C}(t)(l_{C})$ for $T_{P_{C}}(1)$, $P_{C}(t)(l_{C})$ for $T_{P_{C}}(2)$, $P_{C}(t)(l_{C})$ for $T_{P_{C}}(3)$, …$P_{C}(t)(l_{C})$ for $T_{P_{C}}(\infty)$, and $T_{P_{C}}=\{T_{P_{C}}(1),\text{ }T_{P_{C}}(2),\text{ }T_{P_{C}}(3),\text{ }\ldots,\text{ }T_{P_{C}}(\infty)\}$.The cardinality of $f(Z,T)$ transformed problems is countably infinite when we apply $f(Z,T)$ transform on a language that has only finite number of strings in it, e.g., language $L_{C}$.
@@
There is an other decidable language $L_{D}$ and cardinality of set that represents lengths of strings in $L_{D}$ is $\aleph_{0}$, and there exists a Turing machine $P_{D}$ that can recognize $L_{D}$.If $0<T_{P_{D}}\leq\infty$ and $T_{P_{D}}\in\mathbb{Q}^{+}$ then $|T_{P_{D}}|=\aleph_{0}$.We change $P_{D}$ into $P_{D}(t)$ and input of $P_{D}(t)$ is from $L_{D}$.If $|T_{P_{D}}|=\aleph_{0}$ then $f(Z,T)$ transform produces $\aleph_{0}^{\aleph_{0}}$ number of combinations, and each combination is $f(Z,T)$ transformed problem.There are uncountable number of $f(Z,T)$ transformed problems because $2^{\aleph_{0}}\leq\aleph_{0}^{\aleph_{0}}$.
@@
Suppose there is a decidable language $L_{E}$, and we apply $f(Z,T)$ transform on $L_{E}$.If $0<T_{E}\leq\infty$ and $n$ represents the length of strings in $L_{E}$ then Table $1$.represents $f(Z,T)$ transformed problems for language $L_{E}$.If $P_{E}$ is recognizer of $L_{E}$ and $l_{E}\in L_{E}$ then $P_{E}(t)(l_{E})=P_{E}(l_{E})$ for $T_{E}=\infty$ and $L(n\text{ }\infty)=L_{E}$.Therefore, it can be stated that the Language $L_{E}$ is subset of $f(Z,T)$ transformed $L_{E}$, and left most column of of following table represents language $L_{E}$.
@@
$\begin{array}[]{|c|c|c|c|c|c|c|c|c|c|}\hline\cr\mathrm{n}\backslash\mathrm{T_{E}}&1&2&3&4&5&6&7&\ldots&\infty\\ \hline\cr 1&L(1,1)&L(1,2)&L(1,3)&L(1,4)&L(1,5)&L(1,6)&L(1,7)&\ldots&L(1,\infty)\\ \hline\cr 2&{L}(2,1)&L(2,2)&L(2,3)&L(2,4)&L(2,5)&L(2,6)&L(2,7)&\ldots&L(2,\infty)\\ \hline\cr 3&L(3,1)&L(3,2)&L(3,3)&L(3,4)&L(3,5)&L(3,6)&L(3,7)&\ldots&L(3,\infty)\\ \hline\cr 4&L(4,1)&L(4,2)&L(4,3)&L(4,4)&L(4,5)&L(4,6)&L(4,7)&\ldots&L(4,\infty)\\ \hline\cr 5&L(5,1)&L(5,2)&L(5,3)&L(5,4)&L(5,5)&L(5,6)&L(5,7)&\ldots&L(5,\infty)\\ \hline\cr 6&L(6,1)&L(6,2)&L(6,3)&L(6,4)&L(6,5)&L(6,6)&L(6,7)&\ldots&L(6,\infty)\\ \hline\cr 7&L(7,1)&L(7,2)&L(7,3)&L(7,4)&L(7,5)&L(7,6)&L(7,7)&\ldots&L(7,\infty)\\ \hline\cr\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \hline\cr\infty&L(\infty,1)&L(\infty,2)&L(\infty,3)&L(\infty,4)&L(\infty,5)&L(\infty,6)&L(\infty,7)&\ldots&L(\infty,\infty)\\ \hline\cr\end{array}$
Cardinality of Turing machines defined over finite set of states is countably infinite and cardinality of $f(Z,T)$ transformed problems is uncountable, and this difference of cardinalities implies the existence of $f(Z,T)$ transformed problems that are uncomputable by Turing machines defined over finite set of states ( i.e., finite number of tapes and finite set of alphabet).
Through $f(Z,T)$ transform, we can demonstrate the existence of problems that can force different variants of Turing machines (that are defined over finite set of states) to exhibit, at infinitesimal scale, their differential computational characteristics.This work is strictly limited to deterministic variants of Turing machines and computability of $f(Z,T)$ transformed problems; and in future, it will be extended towards nondeterministic Turing machines and complexity theory.
6 Mathematical Formulation of Stipulation Machine Interactions
During construction of $f(Z,T)$ transform we assumed that $W_{P}=1-2^{-T_{P}}$ in (3). But if we assume that $W_{P}=1-2^{-\nicefrac{{T_{P}}}{{c}}}$ and $c$ is some positive constant then we get (34 and 35) after replacing $1-2^{-T_{P}}$ with $1-2^{-\nicefrac{{T_{P}}}{{c}}}$ in (6 and 29).

### v3

@@ 3 Normalization of $f(Z,T)$ Transform
Suppose there is problem $P$ that is computable with input $\psi$, and $\psi\in\Psi$ and $L(P)=\Psi$.We apply $f(Z,T)$ transform on $P(\psi)$ and get $P(t)(\psi)$.Normalization of the transformed problem $P(t)(\psi)$ is defined in following way.
@@
Because $P(\psi)$ is computable problem and $0<T_{P}\leq\infty$ therefore $0\leq f_{n}(P,T)<\infty$.The problem $P(t)(\psi)$ is computable if $10\leq f_{n}(P,T)<$ 1 because $P(t)(\psi)$ halts before Stipulation machine rewrites input. If $N_{P}T_{t}>T_{P}$ then $f_{n}(P,T)>1$ i.e., input of $P(t)$ varies before $P(t)(\psi)$ gets computed and $P(t)$ halts.According to postulate V, no two consecutive inputs of $P(t)$ can identical in a computation.If $f_{n}(P,T)>1$ then Stipulation machine rewrites input on $P(t)$ before $P(t)$ decides and halts on input $\psi$.Consequently, $P(t)$ carries out computation on some parts of each recurring input in some order.Those some parts and some order make arbitrary string $\psi_{f}$ of some arbitrary length, and composition of $\psi_{f}$ depends on $T_{P}$.The arbitrary string $\psi_{f}$ may not be in set $\Psi$.Computation of $P(t)$ on an arbitrary string $\psi_{f}$ is uncomputable.Therefore, if $f_{n}(P,T)>1$ then $P(t)(\psi)$ is not computable but if $10\leq f_{n}(P,T)<$ 1 then $P(t)(\psi)$ is computable.Normalization line for $P(t)(\psi)$ is shown in Fig.3.
@@
Point of inflection and point of convergence are two important points on normalization line. If $f_{n}(P,T)=0$ then transformed problem $P(t)(\psi)$ exist at point of convergence and $f(P,T)=f(P)$ when $f_{n}(P,T)=0$.If $f_{n}(P,T)=1$ then transformed problem $P(t)(\psi)$ exist at point of inflection.The point where $f_{n}(P,T)=1$ is very critical point for computability of problem $P(t)(\psi)$.A small change can make $P(t)(\psi)$ move in-between computable-uncomputable when $f_{n}(P,T)$ exists around point of inflection.If $P(\psi)$ requires $N_{P}$ steps and $P(t)(\psi)$ exists around point of inflection, then slight change in $N_{P}$ (e.g., $N_{P}-n$ or $N_{P}+n$ for some $n\geq 1$) or $T_{P}$ can change the computability of $P(t)(\psi)$.In computation of $P(t)(\psi)$, if $f_{n}(P,T)<1$ then Turing machine carries out computation on input string $(\psi)$ but if $f_{n}(P,T)>1$ then Turing machine carries out computation on some arbitrary string that is composed of some input strings from $\Psi$.We can also represent normalization line of $P(t)(\psi)$ in following way.
@@
The $f(Z,T)$ transformed problems that exist just before or immediately after point of inflection are very helpful in analysing the computational characteristics of different variants of Turing machines.If cardinality of all $f(Z,T)$ transformed problems is higher than cardinality of Turing machines then there must exist $f(Z,T)$ transformed problems that are uncomputable.
4 Cardinality of $f(Z,T)$ Transformed Problems
@@
The $f(Z,T)$ transform is applied to only computable function (or decidable languages) and it produces transformed problems.Suppose there is a language $L_{C}$ and it has only finite number of strings in it, and $l_{C}\in L_{C}$.There is a Turing machine $P_{C}$ that can decide $L_{C}$, i.e., $L(P_{C})=L_{C}$ and $P_{C}$ halts on all inputs.We want to know the cardinality of $f(Z,T)$ transformed problems if we apply $f(Z,T)$ transform on $L_{C}$.We change $P_{C}$ into $P_{C}(t)$ and input is from $L_{C}$ through stipulation machine. If $0<T_{P_{C}}\leq\infty$ and then $|T_{P_{C}}|=\aleph_{0}$, and if $|T_{P_{C}}|=\aleph_{0}$ then $f(Z,T)$ transform produces infinite number of transformed problems in the form of $P_{C}(t)(l_{C})$ for $T_{P_{C}}(1)$, $P_{C}(t)(l_{C})$ for $T_{P_{C}}(2)$, $P_{C}(t)(l_{C})$ for $T_{P_{C}}(3)$, $P_{C}(t)(l_{C})$ for $T_{P_{C}}(4)$, …$P_{C}(t)(l_{C})$ for $T_{P_{C}}(\infty)$, and $T_{P_{C}}=\{T_{P_{C}}(1),\text{ }T_{P_{C}}(2),\text{ }T_{P_{C}}(3),\text{ }\ldots,\text{ }T_{P_{C}}(\infty)\}$.The cardinality of $f(Z,T)$ transformed problems is countably infinite when we apply $f(Z,T)$ transform on a decidable language that has only finite number of strings in it, e.g., language $L_{C}$.
@@
Suppose there is a language $L_{D}$ and cardinality of $L_{D}$ is $\aleph_{0}$, and $l_{D}\in L_{D}$.There exist a Turing machine $P_{D}$ that can decide $L_{D}$, i.e., and $L(P_{D})=L_{D}$ and $P_{D}$ halts on all inputs.We want to know the cardinality of $f(Z,T)$ transformed problems if we apply $f(Z,T)$ transform on $L_{D}$.We change $P_{D}$ into $P_{D}(t)$ and input is from $L_{D}$ through stipulation machine.If $|T_{P_{D}}|=\aleph_{0}$ then $f(Z,T)$ transform produces $\aleph_{0}^{\aleph_{0}}$ number of combinations, and each combination is $f(Z,T)$ transformed language.There are uncountable number of $f(Z,T)$ transformed languages.
@@
Suppose there is a language $L_{E}$ and cardinality of $L_{E}$ is $\aleph_{0}$, and we apply $f(Z,T)$ transform on $L_{E}$.If $n$ represents string number in $L_{E}$ and $0<T_{E}\leq\infty$ then Table $1$.represents $f(Z,T)$ transformed $L_{E}$.If $P_{E}$ is decider of $L_{E}$ and $l_{E}\in L_{E}$ then $P_{E}(t)(l_{E})=P_{E}(l_{E})$ for $T_{E}=\infty$ and $l_{E}(n\text{ }\infty)=L_{E}$.Therefore, it can be stated that the Language $L_{E}$ is subset of $f(Z,T)$ transformed $L_{E}$, and left most column of of following table represents language $L_{E}$.
@@
$\begin{array}[]{|c|c|c|c|c|c|c|c|c|c|}\hline\cr\mathrm{n}\backslash\mathrm{T_{E}}&1&2&3&4&5&6&7&\ldots&\infty\\ \hline\cr 1&l_{E}(1,1)&l_{E}(1,2)&l_{E}(1,3)&l_{E}(1,4)&l_{E}(1,5)&l_{E}(1,6)&l_{E}(1,7)&\ldots&l_{E}(1,\infty)\\ \hline\cr 2&l_{E}(2,1)&l_{E}(2,2)&l_{E}(2,3)&l_{E}(2,4)&l_{E}(2,5)&l_{E}(2,6)&l_{E}(2,7)&\ldots&l_{E}(2,\infty)\\ \hline\cr 3&l_{E}(3,1)&l_{E}(3,2)&l_{E}(3,3)&l_{E}(3,4)&l_{E}(3,5)&l_{E}(3,6)&l_{E}(3,7)&\ldots&l_{E}(3,\infty)\\ \hline\cr 4&l_{E}(4,1)&l_{E}(4,2)&l_{E}(4,3)&l_{E}(4,4)&l_{E}(4,5)&l_{E}(4,6)&l_{E}(4,7)&\ldots&l_{E}(4,\infty)\\ \hline\cr 5&l_{E}(5,1)&l_{E}(5,2)&l_{E}(5,3)&l_{E}(5,4)&l_{E}(5,5)&l_{E}(5,6)&l_{E}(5,7)&\ldots&l_{E}(5,\infty)\\ \hline\cr 6&l_{E}(6,1)&l_{E}(6,2)&l_{E}(6,3)&l_{E}(6,4)&l_{E}(6,5)&l_{E}(6,6)&l_{E}(6,7)&\ldots&l_{E}(6,\infty)\\ \hline\cr 7&l_{E}(7,1)&l_{E}(7,2)&l_{E}(7,3)&l_{E}(7,4)&l_{E}(7,5)&l_{E}(7,6)&l_{E}(7,7)&\ldots&l_{E}(7,\infty)\\ \hline\cr\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \hline\cr\infty&l_{E}(\infty,1)&l_{E}(\infty,2)&l_{E}(\infty,3)&l_{E}(\infty,4)&l_{E}(\infty,5)&l_{E}(\infty,6)&l_{E}(\infty,7)&\ldots&l_{E}(\infty,\infty)\\ \hline\cr\end{array}$
Cardinality of Turing machines is countably infinite and cardinality of $f(Z,T)$ transformed languages is uncountable, and this difference of cardinalities implies the existence of $f(Z,T)$ transformed problems that are uncomputable by Turing machines.
5 Mathematical Formulations of $f(Z,T)$ Transform
During construction of $f(Z,T)$ transform we assumed that $W_{P}=1-2^{-T_{P}}$ in (3). But if we assume that $W_{P}=1-2^{-\nicefrac{{T_{P}}}{{c}}}$ and $c$ is some positive constant then we get (34 and 35) after replacing $1-2^{-T_{P}}$ with $1-2^{-\nicefrac{{T_{P}}}{{c}}}$ in (6 and 32).